OpenStudy (anonymous):

the limit as x goes to 0 of (x^(-1))/((ln(x))^2)

OpenStudy (anonymous):

0

OpenStudy (anonymous):

\[\frac{1}{x (ln{x})^2} \]This limit DNE I think :-)

OpenStudy (anonymous):

Can't be 0...

OpenStudy (anonymous):

It approaches -infinity from 0-, +infinity from 0+

OpenStudy (anonymous):

L'hopitals doesn't apply, does it? I can't remember if you're allowed to use L'hopital's in the case of \[\frac{\infty}{-\infty}\]

OpenStudy (anonymous):

\[ \lim_{x\to0^-}\dfrac{1}{x(\ln(x))^2}=-\infty\\ \lim_{x\to0^+}\dfrac{1}{x(\ln(x))^2}=\infty \]

OpenStudy (anonymous):

Ah yes. What he said.

OpenStudy (anonymous):

It does not, right? The limit is of the form 1/infinity.

OpenStudy (anonymous):

Typo 1/0*

OpenStudy (anonymous):

sorry my mistake, the anwer is infinity, just misread 0 for infinity

OpenStudy (anonymous):

The answer isn't infinity either, the two-sided limit doesn't exist.

OpenStudy (anonymous):

Wait... I'm not sure I agree with the left limit. Does it have a left limit? What happens when you try to evaluate that logarithm...

OpenStudy (anonymous):

I did on the complex plane. Still, the limit DNE even if we are working with only R, we don't have the two sided limit, IMO.

OpenStudy (anonymous):

haha this is math, bmp. Your opinion doesn't enter into it.

OpenStudy (anonymous):

I agree though. I'm inclined to say DNE.

OpenStudy (anonymous):

I was stating my opinion for the reason. That might not be the explanation for it, anyway, on R. But I don't think this exists.

OpenStudy (anonymous):

Don't mind me. I'm just being overly snarky.

OpenStudy (anonymous):

No problem, mate. You are right, I have to avoid using IMO, it's just an old vice :-)

OpenStudy (anonymous):

Could either of you tell me how you got to that because I know L.H is needed but I don't know how to apply it to this

OpenStudy (anonymous):

Well the only way I see of applying L'Hopital's is like this: First of all, it only makes sense to consider the right hand limit. ln(x) does not exist for x<0. Don't rewrite the problem, but keep it as: \[\frac{\frac{1}{x}}{(\ln{x})^2}\] The reason for this is that it means the limit as x approaches 0 will be of indeterminate form infinity/infinity, which allows me to use L'Hopitals rule. So now I just take the derivative of the top and bottom, which gives: \[\frac{\frac{-1}{x^2}}{2\ln(x)*\frac{1}{x}}\] At this point, I'd simplify that and apply L'Hopital's rule again.

OpenStudy (anonymous):

That simplified gives: \[\frac{\frac{-x}{x^2}}{2\ln(x)}= \frac{\frac{-1}{x}}{2\ln(x)}\] Notice that as x approaches 0, this is of indeterminate form -infinity/-infinity. Apply L'hopital's rule to get: \[\frac{\frac{1}{x^2}}{\frac{2}{x}} = \frac{x}{2x^2} = \frac{1}{2x}\] Which clearly approaches infinity.

OpenStudy (anonymous):

And finally, gasp for air.

OpenStudy (anonymous):

It seems that, doing only for R, the limit approaches infinity then. :-) Well done @SmoothMath

OpenStudy (anonymous):

=D